Show that f(x)=(x^3)+(3x^2)-1 has at least one zero in each of the intervals [0,1],[-1,0] and [-3,-2]


1
Expert's answer
2021-01-26T04:17:38-0500

For the proof, we will use the intermediate value theorem states that if f is a continuous function whose domain contains the interval [a, b], then it takes on any given value between f(a) and f(b) at some point within the interval.

This has two important corollaries:

  1. If a continuous function has values of opposite sign inside an interval, then it has a root in that interval (Bolzano's theorem).
  2. The image of a continuous function over an interval is itself an interval.


In our case,

1) "x\in[0,1]"



"\left\{\begin{array}{l}
f(0)=0^3+3\cdot 0^2-1=-1<0\\[0.3cm]
f(1)=1^3+3\cdot 1^2-1=3>0
\end{array}\right.\longrightarrow\\[0.3cm]
\exists c\in(0,1) : f(c)=0"

2) "x\in[-1,0]"



"\left\{\begin{array}{l}
f(0)=0^3+3\cdot 0^2-1=-1<0\\[0.3cm]
f(-1)=(-1)^3+3\cdot(-1)^2-1=1>0
\end{array}\right.\longrightarrow\\[0.3cm]
\exists c\in(-1,0) : f(c)=0"

3) "x\in[-3,-2]"



"\left\{\begin{array}{l}
f(-3)=(-3)^3+3\cdot(-3)^2-1=-1<0\\[0.3cm]
f(-2)=(-2)^3+3\cdot(-2)^2-1=3>0
\end{array}\right.\longrightarrow\\[0.3cm]
\exists c\in(-3,-2) : f(c)=0"

Q.E.D.